Evaluate the following polynomial expression using Horner’s rule. Assume the value of x 

is 3. 

𝑃(𝑥) = 10 𝑥 power 5 + 6 𝑥 power 4 + 3 𝑥 power 3 − 6 𝑥 power 2 + 8𝑥 + 15  Show all the intermediate steps.


1
Expert's answer
2021-05-07T09:49:59-0400

b5 = 10 x 3 =30

b4 = 6 + 3(30) = 96

b3 = 3 + 3(96) = 291

b2 = 6 + 3(291) =879

b1 = 8+3(879) = 2645

b0 = 15 +3(2645) = 7950

7950